OpenStudy (anonymous):

Choose the correct simplification of (a^3b^2)(ab). Answer a^6b^6 a^3b^2 a^4b^3 a^5b^5

OpenStudy (anonymous):

a^4b^3

OpenStudy (anonymous):

see do...'s answer

OpenStudy (anonymous):

bob06 (:

OpenStudy (anonymous):

all you do is multiply each letter by their counterpart (add the exponents :) )

OpenStudy (anonymous):

remember (x^2)(x^5)=x^(2+7) =x^7 and (x^2)^5 = x^(2*5)=x^10

OpenStudy (anonymous):

hey bridget. didnt recognize u w/ ur new pic :)

OpenStudy (anonymous):

(: